One year ago, Mayank's income was 25% more than the income of Prateek. Prateek got 5% increment in his income while due to poor performance, Mayank's income got decreased by 10%. Now by how much percent the income of Prateek is less than the income of Mayank? ​

Grade:10

1 Answers

Harshit Singh
3 years ago
Dear Student
Let us assume that, 1 year ago income of prateek was Rs.100x .

then,

1 year ago, Mayank income = 25% more than 100x = (125 *100x)/100 = Rs.125x .

Now,

Prateek present income = 5% increase than 100x = (105 * 100x)/100 = Rs.105x

and,

Mayank Present income = 10% decrease than 125x = (90 * 125x)/100 = Rs.112.5x

therefore,

Prateek present income less than Mayanak = 112.5x - 105x = Rs.7.5x .

Hence,

% = (7.5x * 100) / 112.5x =6.67%(Ans.)


Thanks